OpenStudy (anonymous):

Tracey has two empty cube-shaped containers with sides of 5 inches and 7 inches. She fills the smaller container completely with water and then pours all the water from the smaller container into the larger container. How deep, to the nearest tenth of an inch, will the water be in the larger container?

OpenStudy (anonymous):

V1 = 5^3 = 125. V1 also = h[(7)(7)]. So, h = 125/49

OpenStudy (anonymous):

What does the h variable represent?

OpenStudy (anonymous):

height or depth, what you are looking for. Because of the 3 dimensions.